# Phenomenological study of the angle between jet axes in heavy-ion collisions

### v1

@@ IV Summary
In this work, we study the angle between the Standard and the Winner-Take-All jet axes $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ in $p$+$p$ and 0-10% Pb+Pb collisions at $\sqrt{s_{NN}}=5.02~{}\mathrm{TeV}$. We mainly investigate the influences of selection bias and the quark/gluon-jet fraction changes on the modification patterns of the $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution in Pb+Pb collisions. The initial $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution in $p$+$p$ collisions is calculated by the Pythia8 event generator. The in-medium jet propagation in nucleus-nucleus collisions is simulated by the Lbt model, which considers both the elastic and inelastic jet-medium interactions.
Firstly, our theoretical results calculated by the Lbt model show that the $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution in Pb+Pb at $\sqrt{s}=5.02~{}\mathrm{TeV}$ is narrower than that in $p$+$p$, which agrees well with the recently reported ALICE measurements. However, the narrowing of $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ seems to violate the $p_{T}$-broadening nature of the jet quenching effect. This is usually explained by the influence of “selection bias”, which usually biases the jet selection by favoring the candidates with less energy loss in A+A collisions. Utilizing a matching-jet method to track the jet evolution in the QGP to reduce the selection bias in the Monte Carlo simulations, we observe that the $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution becomes broader due to the jet-medium interactions. At the same time, by rescaling the quark/gluon-jet fractions in Pb+Pb collisions to be the same as that in $p$+$p$, we find that the fraction change may not significantly influence the modification pattern of jet $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$. On the other hand, we find that the selected jet sample in A+A collisions has a significantly narrower initial $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution than the $p$+$p$ baseline, and such a biased comparison between $p$+$p$ and A+A conceals the actual jet-broadening effect in the experimental measurements. The investigations presented in this paper will deepen our understanding of the relationship between the actual intra-jet modifications in the QGP and the experimental observations.

### v2

@@ IV Summary
In this work, we study the angle between the Standard and Winner-Take-All jet axes $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ in $p$+$p$ and 0-10% Pb+Pb collisions at $\sqrt{s_{NN}}=5.02~{}\mathrm{TeV}$. We mainly investigate the influences of selection bias and the quark/gluon-jet fraction changes on the modification patterns of the $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution in Pb+Pb collisions. The initial $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution in $p$+$p$ collisions is calculated by the Pythia8 event generator. The in-medium jet propagation in nucleus-nucleus collisions is simulated by the Lbt model, which considers both the elastic and inelastic jet-medium interactions.
Firstly, our theoretical results calculated by the Lbt model show that the $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution in Pb+Pb at $\sqrt{s}=5.02~{}\mathrm{TeV}$ is narrower than that in $p$+$p$, which agrees well with the recently reported ALICE measurements. However, the narrowing of $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ seems to violate the nature of intra-jet broadening due to jet quenching. The phenomenon may be attributed to “selection bias”, which skews the jet selection process by preferentially including candidates that exhibit less energy loss in A+A collisions. Utilizing a matching-jet method to track the jet evolution in the QGP to reduce the selection bias in the Monte Carlo simulations, we observe that the $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution becomes broader due to the jet-medium interactions. At the same time, by rescaling the quark/gluon-jet fractions in Pb+Pb collisions to be the same as that in $p$+$p$, we find that the fraction change may not significantly influence the modification pattern of jet $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$. On the other hand, we find that the selected jet sample in A+A collisions has a significantly narrower initial $\Delta R_{\mathrm{axis}}^{\mathrm{WTA}\text{\textendash}\mathrm{Std}}$ distribution than the $p$+$p$ baseline, and such a biased comparison between $p$+$p$ and A+A might conceal the actual intra-jet broadening effect in the experimental measurements. The investigations presented in this paper will deepen our understanding of the relationship between the actual intra-jet modifications in the QGP and the experimental observations. However, the methodology currently utilized in this study cannot be directly applied to experimental measurements. Recently, we have noticed that Z/$\gamma$-tagged jets or high-$p_{\mathrm{T}}$-hadron-tagged jets can be used to reduce selection bias and can apply to real experimental measurements [95, 97, 96, 101, 102]. Specifically, by using tagger particles rather than jets for event selection, the selection bias is only induced by the choice of tagger. Since tagger particles do not interact with the QGP medium, this approach significantly reduces the impact of selection bias on jet quenching studies. Our future research will employ Z/$\gamma$-tagged jets to minimize selection bias and explore intra-jet broadening phenomena in heavy-ion collisions.